OpenStudy (anonymous):

lim sin h/2h h>0 using squeezing theorem .

hartnn (hartnn):

do u know basic limit identity ? \(\huge\lim \limits_{h \rightarrow0}\frac{sin h}{h}=1\)

hartnn (hartnn):

from your Q, \(\huge\lim \limits_{h \rightarrow0}\frac{sin h}{2h}=\huge \frac{1}{2}\lim \limits_{h \rightarrow0}\frac{sin h}{h}=\frac{1}{2}\) got that ?

OpenStudy (anonymous):

how it can be 1/2 ?

hartnn (hartnn):

the 2 in the denominator is factored out as (1/2) because constants can be taken out of limits. what remains is a standard limit formula which equals 1.

OpenStudy (anonymous):

@hartnn he is asking by squeeze principle..............this is the general way

hartnn (hartnn):

ohh...using squeezing theorem...

OpenStudy (anonymous):

ya...

hartnn (hartnn):

then u need to know this fact : \( \cos x \lt \frac{\sin x}{x} \lt 1\)

hartnn (hartnn):

then using squeeze theorem u can directly say that \(\huge\lim \limits_{h \rightarrow0}\frac{sin h}{h}=1\)

hartnn (hartnn):

or, if u want to include 2 from the beginning, write \(\large\frac{\cos x}{2} \lt \frac{\sin x}{2x} \lt \frac{1}{2}\) then using squeeze, u can direct get your answer.

OpenStudy (anonymous):

thanks ! it's really useful for me .

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

lim 2-cos3x-cos4x/x x>0 using squeezing theorem .

hartnn (hartnn):

u must be knowing \(-1 \lt \cos x \lt 1\) right ?

hartnn (hartnn):

rearrange that to get \(\frac{1-cos x}{x}\) in between